# Two-Player Zero-Sum Games with Bandit Feedback

### v1

@@ 5 ETC-TPZSG-AE Algorithm and Regret Analysis
$\hat{A}(i^{\prime},j)-\varepsilon_{t}\leq\hat{A}(i,j)\leq\hat{A}(i,j^{\prime})+\varepsilon_{t},\forall i^{\prime},\forall j^{\prime}$
@@
18:            Reset $S_{t+1}$ by "Action Pair Elimination" step
Algorithm 2 ETC-TPZSG-AE
Theorem 5.1.
where $\lambda\geq\sqrt{\frac{16\sigma^{2}e}{T}}$ and, $S_{A_{1}}=\{(i,j)\in S_{A}:\Delta_{ij}>\lambda\}$ and $S_{A_{2}}=\{(i,j)\in S_{A}:0<\Delta_{ij}\leq\lambda\}$ are two subsets of the action pairs.
Theorem 5.2.
$R_{T}\leq\sum_{(i,j)\in S_{A_{1}}}\Big{(}\Delta_{ij}+\frac{768\sigma^{2}}{\Delta_{ij}}+\frac{256\sigma^{2}}{\Delta_{ij}}\ln{\Big{(}\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big{)}}\Big{)}+\sum_{(i,j)\in S_{A_{2}}}\Big{(}\frac{512\sigma^{2}}{\lambda}+\lambda T\Big{)}$ (22)
We simplify the regret bound to $O(\frac{\log(T\Delta^{2})}{\Delta})$, as the logarithmic term dominates since by setting $\lambda=\sqrt{\frac{16\sigma^{2}e}{T}}$, we ensure that the term $\lambda T$ is bounded by $\sqrt{16\sigma^{2}eT}$ which is at most $\frac{e}{\Delta_{ij}}$ when $\Delta_{ij}\leq\lambda$. Our results are consistent with the regret bound derived by auer2010ucb . The proofs of the theorems are provided in Appendix C.
In this section, we analyze two distinct regret formulations for Algorithm 2, Nash regret, defined in (3.4), and external regret, defined in (3.3). The first, denoted by $R_{T}^{*}$, evaluates performance against the NE, capturing the cumulative loss relative to the equilibrium strategy. In contrast, the regret analysis by $R_{T}$ allows us to consider the losses of both players with respect to their individual best responses at each round, providing a broader and more flexible notion of regret. By the inequality $\Delta{ij}^{*}\leq\Delta{ij}$ for all action pairs, it follows that $R_{T}^{*}\leq R_{T}$. This outcome aligns with the interpretation that the expected regret against the NE establishes a more specific comparison and thus, it naturally provides a tighter upper bound on the expected regret.
@@ 6 Experiments
In this section, we present a set of simulations to support our theoretical findings and demonstrate the performance of the proposed algorithms. The experiments are conducted with fixed time horizons of $T=10^{3}$ and $T=10^{4}$. Additional details are provided in Appendix D.
(a) The expected regret of ETC-TPZSG with $k$ in (16) and the upper bound in (17)
(b) The cumulative regrets of ETC-TPZSG, ETC-TPZSG-AE and Tsallis-INF from ito2025instance
(c) The theoretical regret bounds for ETC-TPZSG, ETC-TPZSG-AE and Tsallis-INF from ito2025instance
Figure 1: Regret performance comparison of the algorithms
Figure 1(a) shows the expected regret bound, averaged over $10^{5}$ simulation runs, of ETC-TPZSG using $k$ in (16) and the upper bound in (17). We consider a setting with $N=2$, where the row player has two actions, the column player has one, and the first row corresponds to the NE. The suboptimality gap $\Delta$ varies from 0 to 1 and the subgaussian parameter is set to $\sigma=0.5$. The results align with the bound in Theorem (4.1).
In Figure 1(b), we compare ETC-TPZSG, ETC-TPZSG-AE, and Tsallis-INF ito2025instance , the only existing method with instance-dependent bounds for TPZSGs, based on cumulative regret, computed as the absolute difference between the game value and the payoff from the action played to avoid negative values, averaged over $10^{3}$ runs. The results show that ETC-TPZSG-AE achieves lower regret than ETC-TPZSG, highlighting the effectiveness of its action pair elimination strategy. We use a $2\times 2$ game matrix by Gaussian payoffs and $k$ exploration rounds per action pair for ETC-TPZSG is randomly selected from a predefined list. On the other hand, Figure 1(c) compares their theoretical regret bounds ($O(\cdot)$ notation) for $\Delta=0.5$, which aligns with the results in Figure 1(b).
@@ 7 Discussion
We investigate a TPZSG with bandit feedback, where the payoff matrix is unknown and must be learned through player interactions. This setting is challenging, as players must estimate payoffs while making strategic decisions in an adversarial environment. We adopt the ETC algorithm due to its simplicity, widespread use and lack of prior analysis in this context. We also integrate an elimination based method, allowing the systematic removal of suboptimal action pairs, thereby improving convergence to the equilibrium and reducing unnecessary exploration on suboptimal ones. A key contribution of our work is the derivation of instance-dependent upper bounds on the expected regret for both algorithms, which has received limited attention in the literature on zero-sum games.
While our study focuses on pure strategy learning in a zero-sum game with bandit feedback and provides a theoretical expected regret bounds, several directions remain open for future study. One interesting extension is to consider games where the equilibrium is mixed. While the algorithms provided can be used to identify the support of the equilibrium, another approach should be used to efficiently converge to a mixed equilibrium.
Finally, an important and relatively unexplored direction is fairness in zero-sum games. For example, introducing mechanisms that ensure similar action pairs are explored equally can promote fairness in strategy estimation. This type of fair play mechanism could be essential in TPZSG settings, thus it can promote balanced exploration and improve overall strategy estimation. On the other hand, by integrating fairness based algorithms or constraints, it may be possible to generate game environments that ensure balanced opportunities for all players.
@@ Appendix A Subgaussian Properties
This section collects some supplementary results from [20] that are used repeatedly throughout our proofs. These results concern standard properties of subgaussian random variables such as tail inequalities. They are included here for completeness and easy reference since they play a critical role in our analyses. The statements are not original, we include only the parts that are directly useful for our purposes.
Theorem A.1 ([20, Theorem 5.3]).
@@
    $cX$ is $|c|\sigma$-subgaussian for all $c\in\mathbb{R}.$
@@
Appendix B Proof of Theorem 4.1
To analyze the Nash regret of the ETC-TPZSG algorithm, we begin by decomposing the total expected regret into two phases: the exploration phase and the committing phase. During exploration, each action pair is played a fixed number of times $k$ to estimate their mean rewards, potentially incurring regret if suboptimal action pairs are played. Once the algorithm commits to the empirically best action pair, which is the pure NE, the regret accumulates only if this action pair is not the true optimal one. Our goal is to bound the expected regret by quantifying the probability of committing to a suboptimal action pair, which we can call it as a misidentified NE, based on the estimates of their payoffs obtained during exploration phase.
@@
Let us start by considering the first approach. We can write it as
@@
$\leq\sqrt{\mathbb{P}(\hat{A}(i,j^{\prime})-\hat{A}(i^{\prime},j^{\prime})\leq\Delta_{i^{\prime}j^{\prime}}^{\text{max}})\ \mathbb{P}(\hat{A}(i^{\prime},j^{\prime})-\hat{A}(i^{\prime},j)\leq\Delta_{i^{\prime}j^{\prime}}^{\text{min}})}$ (28)
@@
To get (26), we utilize the fact that the probability of the intersection of multiple events is at most that of any individual event, that is, for any collection of events $e_{1},e_{2},\ldots,e_{n}$, we have $\mathbb{P}(e_{1}\cap e_{2}\cap\cdots\cap e_{n})\leq\mathbb{P}(e_{i})$ for all $i\in{1,\ldots,n}$. The inequality (28) holds since $\Delta_{i^{\prime}j^{\prime}}^{\text{max}}\geq 0,\Delta_{i^{\prime}j^{\prime}}^{\text{min}}\geq 0$, which are true from their definitions. For the step (29), we apply Theorem A.1 with $\sqrt{2\sigma^{2}/k}$-subgaussian random variable.
At this point, we need to show that the differences $\hat{A}(i,j)-\hat{A}(x,y)$ are $\sqrt{2\sigma^{2}/k}$-subgaussian where $(i,j)$ and $(x,y)$ are any action pairs in $S_{A}$. To establish this, we utilize Lemma A.1, which provides key properties of subgaussian random variables. Specifically, both $\hat{A}(i,j)$ and $\hat{A}(x,y)$ are $\sqrt{\sigma^{2}/k}$-subgaussian as the average payoffs are computed by the equation (3) and during exploration phase, each action pair is played $k$ times. By the lemma, the difference of two independent subgaussian random variables with parameters $\sqrt{\sigma^{2}/k}$ is subgaussian with parameter $\sqrt{2\sigma^{2}/k}$, which follows that $\hat{A}(i,j)-\hat{A}(x,y)$ is $\sqrt{2\sigma^{2}/k}$-subgaussian for any action pairs $(i,j)$ and $(x,y)$.
Then, to get the last inequality (31), we observe that
@@
As we mention before, there is an alternative approach to bound $\mathbb{P}(E_{1}\cap E_{2})$, which means to find a bound for the probability of a misidentified NE. It involves an inequality based on the minimum of the individual probabilities. Similarly, let follow this:
@@
$\leq\min\{\mathbb{P}(\hat{A}(i,j^{\prime})-\hat{A}(i^{\prime},j^{\prime})\leq\Delta_{i^{\prime}j^{\prime}}^{\text{max}}),\mathbb{P}(\hat{A}(i^{\prime}j^{\prime})-\hat{A}(i^{\prime}j)\leq\Delta_{i^{\prime}j^{\prime}}^{\text{min}})\}$ (38)
@@
where $\mathbb{E}[n_{ij,T}]$ is the expected number of playing the action pair $(i,j)$. We already know that each action pair $(i,j)\in S_{A}$ is played at least $k$ times for exploration. After the exploration step, we will consider playing the misidentified NE $(i^{\prime},j^{\prime})$ instead of real one $(i^{*},j^{*})$ because we expect the players to select NE action pair as optimal after exploration phase. Since we already have the probability of playing a misidentified NE from (33) or (43), we can write
@@
Hence, the proof is concluded.
$\Box$
Appendix C Proof of Theorem 5.1
According to the algorithm, $\sqrt{\frac{16\sigma^{2}e}{T}}$ is a critical value for $\hat{\Delta}_{t}$ since it takes this value in the last round. In this way, let define a threshold parameter as $\lambda\geq\sqrt{\frac{16\sigma^{2}e}{T}}$, which gives us two subsets of actions pairs such that $S_{A_{1}}=\{(i,j)\in S_{A}:\Delta_{ij}>\lambda\}$ and $S_{A_{2}}=\{(i,j)\in S_{A}:0<\Delta_{ij}\leq\lambda\}$. Since $\Delta_{ij}^{*}\leq\Delta_{ij}$ ensures the bounds are preserved, it makes sense to use these action pair sets in the analysis. Thus, we can write the Nash regret as
@@
To analyze the Nash regret of the ETC-TPZSG-AE algorithm, it is essential to consider various scenarios that may lead to suboptimal outcomes. Each of these scenarios contributes to the total Nash regret, as they either involve the incorrect elimination of the optimal action pair or the unnecessary selection of suboptimal pairs. The former results in discarding the optimal strategy because of insufficient evidence, while the latter leads to spend more time with action pairs that are unlikely to be part of any near optimal equilibrium.
The algorithm employs a scheduling mechanism for $\Delta_{ij}$, where $\hat{\Delta}_{t}$ is halved in each round. This approach is necessary because the true values of $\Delta{ij}$ are unknown to the players. We use $\hat{\Delta}_{t}$ to determine both the exploration time and the tolerance level in the $\varepsilon$-NE property. Specifically, when $\hat{\Delta}_{t}$ is large, the algorithm plays each action pair fewer times and keeps more pairs during the elimination step, since the threshold for elimination is relatively loose. As $\hat{\Delta}_{t}$ decreases, the algorithm allocates more exploration to action pairs that are closer to being part of a NE, because those that are clearly suboptimal have already been eliminated. This adaptive process enables more precise identification of near-optimal strategies by $\varepsilon$-NE condition while minimizing regret.
@@
Let explain briefly in which situations the algorithm eliminates or keep an action pair. For action pair $(i,j)$, if $\varepsilon$-NE property does not hold in round $t=t_{ij}$ which implies that at least one of the inequalities fails, then it will be eliminated in round $t_{ij}$.
Furthermore, we have $\varepsilon_{t_{ij}}=\sqrt{\frac{4\sigma^{2}}{k_{t_{ij}}}\ln{\Big{(}\frac{\hat{\Delta}_{t_{ij}}^{2}T}{16\sigma^{2}}\Big{)}}}\leq\frac{\hat{\Delta}_{t_{ij}}}{2}=\hat{\Delta}_{t_{ij}+1}<\frac{\Delta_{ij}}{4}$. In other words, if there exist $i^{\prime}\in S_{x}$ or $j^{\prime}\in S_{y}$, such that one of the following is true:
@@
where $\varepsilon_{t}=\sqrt{\frac{4\sigma^{2}}{k_{t}}\ln{\Big{(}\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\Big{)}}}.$ To derive inequality (48), we use the fact that the probability of the intersection of multiple events is always less than or equal to the probability of any of individual events. Specifically, since we have $m$ actions in $S_{x}$, for any collection of events $e_{1},e_{2},\cdots,e_{m}$, it holds that $\mathbb{P}(e_{1}\cap e_{2}\cap\cdots\cap e_{m})\leq\mathbb{P}(e_{i})$ for all $i=1,2,\cdots,m$. In the last inequality (49), we apply Theorem A.1 with $\sqrt{2\sigma^{2}/k_{t}}$-subgaussian random variable. The subgaussian parameter is derived by using Lemma A.1 and we note that each action pair is played $k_{t}$ times.
@@
On the other hand, we note that there is no need to consider the action pairs $(i,j)\in S_{A_{2}}$ because we handle the elimination of suboptimal pairs before the final round. Thus, using (45), each action pair $(i,j)$ is played at most $k_{t_{ij}}$ times such that $k_{t_{ij}}=\Big{\lceil}\frac{16\sigma^{2}}{\hat{\Delta}_{t_{ij}}^{2}}\ln{(\frac{\hat{\Delta}_{t_{ij}}^{2}T}{16\sigma^{2}})}\Big{\rceil}\leq\Big{\lceil}\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big{(}\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big{)}}\Big{\rceil}$.
@@
If the algorithm keeps a suboptimal action pair $(i,j)$ up to round $t_{*}$, the events specified in (46) must hold with parameter $\varepsilon_{t_{*}}$. Consequently, by applying (52), the probability of keeping a suboptimal action pair can be bounded by $\frac{4N\sigma^{2}}{\hat{\Delta}_{t_{*}}^{2}T}$. Thus, we can write the regret contribution of this case as
@@
Case C.4.
A suboptimal action pair $(i,j)$ in the set $S_{A_{2}}$ remains in the game as a unique action pair in $S_{t_{ij}}$.
@@
This regret contribution accounts for scenarios where the algorithm may eliminate all action pairs except one. If the remaining action pair is suboptimal, the algorithm will continue to play with this action pair for the remaining rounds. Since the action pair selection persists up to time step $T$, resulting in the additional term in the regret bound.

### v2

@@ 6 ETC-TPZSG-AE-NUE Algorithm
$\hskip 35.56593pt\hat{A}(i^{\prime},j)-\varepsilon_{t}\leq\hat{A}(i,j)\leq\hat{A}(i,j^{\prime})+\varepsilon_{t},\forall i^{\prime},\forall j^{\prime}$
@@
18:     $\hat{P}_{ij,t}$ by solving the linear system in (24).
19:     $S_{t+1}$ by "Action Pair Elimination" step
20:     $\hat{\Delta}_{t+1}=\frac{\hat{\Delta}_{t}}{2}$ where $\hat{\Delta}_{t}=2^{-t+2}\sigma$
Algorithm 3 ETC-TPZSG-AE-NUE
Theorem 4.
The Nash regret of Algorithm 3 is upper bounded by
$R_{T}^{*}\leq\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}\Big(1+\frac{768\sigma^{2}}{\Delta_{ij}^{2}}+\frac{512\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big)+\sum_{(i,j)\in S_{A_{2}}}\Big(\Delta_{ij}^{*}\frac{512\sigma^{2}}{\lambda^{2}}+\Delta_{ij}^{*}T\Big)$ (26)
where $\lambda\geq 4\sigma\sqrt[4]{e/T}$ and, $S_{A_{1}}=\{(i,j)\in S_{A}:\Delta_{ij}>\lambda\}$ and $S_{A_{2}}=\{(i,j)\in S_{A}:0<\Delta_{ij}\leq\lambda\}$ are two subsets of the action pairs.
Proof Sketch.
Having the same structure as Algorithm 2 and ensuring that each action pair is explored at least $k_{t}$ times enables us to use the same proof steps; however, this modification affects the bound on the number of plays when the algorithm eliminates a suboptimal action pair in round $t_{ij}$. Detailed explanation is given in Appendix D. We note that since we have larger $\lambda$ here, it makes this bound tighter compared to the results of Algorithm 2. ∎
In conclusion, Algorithm 3 aims that the strategies of players converge to the corresponding pure strategies over time, which allows more focused exploration on actions near the pure NE and thus enables to eliminate suboptimal action pairs more effectively. In addition, the regret analysis of Algorithm 3 provides a tighter upper bound, highlighting the improved efficiency of the algorithm in learning pure NE.
@@ 7 Experiments
In this section, we present a set of simulations to support our theoretical findings and demonstrate the performance of the proposed algorithms. The experiments are conducted with a fixed time horizon $T=10^{3}$ for Figure 1 and $T=10^{4}$ for other simulations. Additional details are provided in Appendix E.
Figure 1: The expected regret of ETC-TPZSG with $k$ in (17) and the upper bound in (18)
Figure 2: The cumulative regrets of proposed algorithms and Tsallis-INF with large $\Delta$
Figure 3: The cumulative regrets of proposed algorithms and Tsallis-INF with small $\Delta$
Figure 1 shows the expected regret, averaged over $10^{5}$ runs, of ETC-TPZSG using $k$ in (17) and the upper bound in (18). We consider a setting with $N=2$, where the row player has two actions, the first corresponds to NE with mean zero and the second with $\Delta$, and the column player has one action. The suboptimality gap $\Delta$ varies from 0 to 1 and $\sigma=0.5$. The fluctuation around $\Delta=0.1$ arises since, up to this point, the exploration time $k$ equals one, after which it increases suddenly. The results align with the bound in Theorem 1.
In Figures 3 and 3, we compare the proposed algorithms and Tsallis-INF [18], the only existing method with instance-dependent bounds for TPZSGs, based on cumulative regret, computed as the absolute difference between the game value and the payoff from the action played to avoid negative values, averaged over $10^{2}$ runs. We use a $2\times 2$ game matrix by Gaussian payoffs with variance $\sigma^{2}$ and $k$ exploration rounds per action pair for ETC-TPZSG is randomly selected from a predefined list. We note that a smaller $\sigma$ leads to a smaller $\Delta$.
We evaluate the algorithms using the payoff matrices with large and small $\Delta$ in Figures 3 and 3, respectively, to examine how $\Delta$ affects the performance of the algorithms. This comparison highlights the importance of instance-dependent regret analysis. The results show that the algorithms incorporating elimination methods achieve lower regret than the ETC-TPZSG and Tsallis-INF algorithms, highlighting the effectiveness of adaptive elimination and exploration.
On the other hand, when $\Delta$ is small, the Tsallis-INF algorithm fails to correctly identify pure NE as we see an increase in its cumulative regret in Figure 3. In contrast, Algorithms 2 and 3 perform better. Thus, we also emphasize that our algorithms preserve the pure NE property, which partly explains why they perform better than Tsallis-INF in games that admit pure NE. Experimental details and further results with different game matrices are provided in Appendix E.
@@ 8 Discussion
We investigate a TPZSG setting with bandit feedback, where the true payoff matrix is unknown and must be learned through player interactions. This setting is challenging, as players must estimate payoffs while making strategic decisions in an adversarial environment. We adopt the ETC algorithm due to its simplicity, widespread use and lack of prior analysis in this context. We integrate it to an adaptive elimination method, allowing the systematic removal of suboptimal action pairs, thereby improving convergence to the equilibrium and reducing unnecessary exploration on suboptimal ones. Finally, we combine the adaptive elimination algorithm with a non-uniform exploration approach to explore action pairs more effectively. A key contribution of our work is the derivation of instance-dependent upper bounds on the expected regret for proposed algorithms, which has received limited attention in the literature on zero-sum games.
Although our study focuses on pure strategy learning in a zero-sum game with bandit feedback and provides instance-dependent regret bounds, several directions remain open for future study. An interesting extension is to explore games with a mixed strategy equilibrium, which is more challenging because of the convergence dynamics to true strategies, and also more complex to analyze regret compared to a pure strategy setting. Another direction is to analyze general-sum games, where the player interactions are not strictly adversarial, which requires to consider additional factors to learn optimal strategies.
Finally, an important and relatively unexplored direction is fairness in zero-sum games. For example, introducing mechanisms that ensure similar action pairs are explored equally can promote fairness in strategy estimation. This type of fair play mechanism could be essential in TPZSG settings, thus it can promote balanced exploration and improve overall strategy estimation. On the other hand, by integrating fairness based algorithms or constraints, it may be possible to generate game environments that ensure equal opportunities for all players.
Acknowledgments.
The authors would like to thank Debabrota Basu, Thomas Kleine Buening and Ronald Ortner for fruitful discussions on this topic.
@@ Appendix A Subgaussian Properties
This section collects some supplementary results from [20] that are used repeatedly throughout our proofs. These results concern standard properties of subgaussian random variables such as tail inequalities. They are included here for completeness and easy reference as they play a critical role in our analyses. The statements are not original, we include only the parts that are directly useful for our purposes.
Theorem 5 ([20, Theorem 5.3]).
@@
    cXcX is $|c|\sigma$-subgaussian for all $c\in\mathbb{R}.$
@@
Appendix B Proof of Theorem 1
To analyze the Nash regret of the Algorithm 1, we begin by decomposing the total expected regret into two phases: the exploration phase and the committing phase. During exploration, each action pair is played a fixed number of times $k$ to estimate their mean rewards, potentially incurring regret if suboptimal action pairs are played. Once the algorithm commits to the empirically best action pair, which is the pure NE, the regret accumulates only if this action pair is not the true optimal one. Our goal is to bound the expected regret by quantifying the probability of committing to a suboptimal action pair, which we can call it as a misidentified NE, based on the estimates of their payoffs obtained during the exploration phase.
@@ Appendix B Proof of Theorem 1
Let us start by considering the first approach. We can write it as follows:
@@
$\leq\sqrt{\mathbb{P}(\hat{A}(i,j^{\prime})+A(i,j^{\prime})-\hat{A}(i^{\prime},j^{\prime})-A(i^{\prime},j^{\prime})\leq\Delta_{i^{\prime}j^{\prime}}^{\text{max}})}$
$\quad\times\sqrt{\mathbb{P}(\hat{A}(i^{\prime},j^{\prime})+A(i^{\prime},j^{\prime})-\hat{A}(i^{\prime},j)-A(i^{\prime},j)\leq\Delta_{i^{\prime}j^{\prime}}^{\text{min}})}$ (32)
@@
To get (30), we utilize the fact that the probability of the intersection of multiple events is at most that of any individual event, that is, for any collection of events $e_{1},e_{2},\ldots,e_{n}$, we have $\mathbb{P}(e_{1}\cap e_{2}\cap\cdots\cap e_{n})\leq\mathbb{P}(e_{i})$ for all $i\in{1,\ldots,n}$. The inequality (32) holds by using the followings:
$\mathbb{P}(\hat{A}(i,j^{\prime})-\hat{A}(i^{\prime},j^{\prime})\leq 0)=\mathbb{P}(\hat{A}(i,j^{\prime})+A(i,j^{\prime})-\hat{A}(i^{\prime},j^{\prime})-A(i^{\prime},j^{\prime})\leq A(i,j^{\prime})-A(i^{\prime},j^{\prime}))$ (38)
$\mathbb{P}(\hat{A}(i^{\prime},j^{\prime})-\hat{A}(i^{\prime},j)\leq 0)=\mathbb{P}(\hat{A}(i^{\prime},j^{\prime})+A(i^{\prime},j^{\prime})-\hat{A}(i^{\prime},j)-A(i^{\prime},j)\leq A(i^{\prime},j^{\prime})-A(i^{\prime},j))$ (39)
Then, we apply $\Delta_{i^{\prime}j^{\prime}}^{\text{max}}\geq A(i,j^{\prime})-A(i^{\prime},j^{\prime}),\Delta_{i^{\prime}j^{\prime}}^{\text{min}}\geq A(i^{\prime},j^{\prime})-A(i^{\prime},j)$, which are true from their definitions. For step (33), we apply Theorem 5 with $\sqrt{2\sigma^{2}/k}$-subgaussian random variable. At this point, we need to show that the differences $\hat{A}(i,j^{\prime})+A(i,j^{\prime})-\hat{A}(i^{\prime},j^{\prime})-A(i^{\prime},j^{\prime})$ and $\hat{A}(i^{\prime},j^{\prime})+A(i^{\prime},j^{\prime})-\hat{A}(i^{\prime},j)-A(i^{\prime},j)$ are $\sqrt{2\sigma^{2}/k}$-subgaussian but it is enough to check that for any action pairs $(i,j)$ and $(x,y)$ in $S_{A}$, $\hat{A}(i,j)-\hat{A}(x,y)$ are $\sqrt{2\sigma^{2}/k}$-subgaussian. To establish this, we utilize Lemma 1, which provides key properties of subgaussian random variables. Specifically, both $\hat{A}(i,j)$ and $\hat{A}(x,y)$ are $\sqrt{\sigma^{2}/k}$-subgaussian as the average payoffs are computed by the equation (4) and during the exploration phase, each action pair is played $k$ times. By the lemma, the difference of two independent subgaussian random variables with parameters $\sqrt{\sigma^{2}/k}$ is subgaussian with parameter $\sqrt{2\sigma^{2}/k}$, which follows that $\hat{A}(i,j)-\hat{A}(x,y)$ is $\sqrt{2\sigma^{2}/k}$-subgaussian for any action pairs $(i,j)$ and $(x,y)$. To get the last inequality (35), we observe that
@@
As we mention before, there is an alternative approach to bound $\mathbb{P}(E_{1}\cap E_{2})$ to find a bound for the probability of a misidentified NE. It involves an inequality based on the minimum of individual probabilities. Similarly, let follow this:
@@
where $\mathbb{E}[n_{ij,T}]$ is the expected number of playing the action pair $(i,j)$. We already know that each action pair $(i,j)\in S_{A}$ is played at least $k$ times for exploration. After the exploration step, we will consider playing the misidentified NE $(i^{\prime},j^{\prime})$ instead of true one $(i^{*},j^{*})$ because we expect the players to select NE action pair as optimal after the exploration phase. Since we already have the probability of playing a misidentified NE from (37) or (48), we can write
@@
Hence, the proof is concluded. $\Box$
Appendix C Proof of Theorem 2
According to the algorithm, $4\sigma\sqrt{e/T}$ is a critical value for $\hat{\Delta}_{t}$ since it takes this value in round $\frac{1}{2}\log_{2}\frac{T}{e}$. In this way, let define a threshold parameter as $\lambda\geq 4\sigma\sqrt{e/T}$, which gives us two subsets of actions pairs such that $S_{A_{1}}=\{(i,j)\in S_{A}:\Delta_{ij}>\lambda\}$ and $S_{A_{2}}=\{(i,j)\in S_{A}:0<\Delta_{ij}\leq\lambda\}$. Since $\Delta_{ij}^{*}\leq\Delta_{ij}$ ensures the bounds are preserved, it makes sense to use these action pair sets in the analysis. Thus, we can write the Nash regret as
@@ Appendix C Proof of Theorem 2
To analyze the Nash regret of Algorithm 2, it is essential to consider various scenarios that may lead to suboptimal outcomes. Each of these scenarios contributes to the total Nash regret, as they either involve the incorrect elimination of the optimal action pair or the unnecessary selection of suboptimal pairs. The former results in discarding the optimal strategy because of insufficient evidence, while the latter leads to spend more time with action pairs that are unlikely to be part of any near optimal equilibrium.
The algorithm employs a scheduling mechanism for $\Delta_{ij}$, where $\hat{\Delta}_{t}$ is halved in each round. This approach is necessary because the true values of $\Delta_{ij}$ are unknown to the players. We use $\hat{\Delta}_{t}$ to determine both the exploration time and the tolerance level in the $\varepsilon$-NE property. Specifically, when $\hat{\Delta}_{t}$ is large, the algorithm plays each action pair fewer times and keeps more pairs during the elimination step, since the threshold for elimination is relatively loose. As $\hat{\Delta}_{t}$ decreases, the algorithm allocates more exploration to action pairs that are closer to being part of a NE, because those that are clearly suboptimal have already been eliminated. This adaptive process enables a more precise identification of near-optimal strategies by $\varepsilon$-NE condition while minimizing regret.
@@
Let explain briefly in which situations the algorithm eliminates or keeps an action pair. For action pair $(i,j)$, if $\varepsilon$-NE property does not hold in round $t=t_{ij}$ which implies that at least one of the inequalities fails, then it will be eliminated in round $t_{ij}$.
Furthermore, we have $\varepsilon_{t_{ij}}=\sqrt{\frac{4\sigma^{2}}{k_{t_{ij}}}\ln{\Big(\frac{\hat{\Delta}_{t_{ij}}^{2}T}{16\sigma^{2}}\Big)}}\leq\frac{\hat{\Delta}_{t_{ij}}}{2}=\hat{\Delta}_{t_{ij}+1}<\frac{\Delta_{ij}}{4}$. In other words, if there exists $i^{\prime}\in S_{x}$ or $j^{\prime}\in S_{y}$, such that one of the following is true:
@@
where $\varepsilon_{t}=\sqrt{\frac{4\sigma^{2}}{k_{t}}\ln{\Big(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\Big)}}.$ To derive inequality (53), we use the fact that the probability of the intersection of multiple events is always less than or equal to the probability of any of individual events. Specifically, since we have $m$ actions in $S_{x}$, for any collection of events $e_{1},e_{2},\cdots,e_{m}$, it holds that $\mathbb{P}(e_{1}\cap e_{2}\cap\cdots\cap e_{m})\leq\mathbb{P}(e_{i})$ for all $i=1,2,\cdots,m$. In the last inequality (54), we apply Theorem 5 with $\sqrt{2\sigma^{2}/k_{t}}$-subgaussian random variable. The subgaussian parameter is derived by using Lemma 1 and we note that each action pair is played $k_{t}$ times.
@@
On the other hand, we note that there is no need to consider the action pairs $(i,j)\in S_{A_{2}}$ because we handle the elimination of suboptimal pairs before the final round. Thus, using (50), each action pair $(i,j)$ is played at most $k_{t_{ij}}$ times such that
$k_{t_{ij}}=\Big\lceil\frac{16\sigma^{2}}{\hat{\Delta}_{t_{ij}}^{2}}\ln{(\frac{\hat{\Delta}_{t_{ij}}^{2}T}{16\sigma^{2}})}\Big\rceil\leq\Big\lceil\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big\rceil.$
@@
If the algorithm keeps a suboptimal action pair $(i,j)$ up to round $t_{*}$, the events specified in (51) must hold with parameter $\varepsilon_{t_{*}}$. Consequently, by applying (57), the probability of keeping a suboptimal action pair can be bounded by $\frac{16\sigma^{2}}{\hat{\Delta}_{t_{*}}^{2}T}$. Thus, we can write the regret contribution of this case as
@@
Case 4.
A suboptimal action pair $(i,j)$ from the set $S_{A_{2}}$ is selected as the unique action pair in $S_{t_{ij}}$ for remaining rounds.
@@
This regret contribution accounts for scenarios where the algorithm selects suboptimal action pair as the pure NE. In other words, if the selected action pair is suboptimal, the algorithm will continue to play with this action pair for the remaining rounds. Since the action pair selection persists up to time step $T$, resulting in an additional term in the regret bound.